OpenStudy (helpjebal):

Find solutions for the equation sin^2 x + sin x = 0

OpenStudy (helpjebal):

I've gotten to: sin x(sin x + 1) = 0 not sure where to go from here...

Nnesha (nnesha):

set it equal to zero example \[\rm x(y+z)=0~~~\rightarrow ~~x=0~~~~and~~y+z=0\]

OpenStudy (helpjebal):

okay so sin x = 0 sin x + 1 = 0

Nnesha (nnesha):

right now solve sinx+1=0 for sinx

OpenStudy (helpjebal):

sin x +1 = 0 sin x = -1

Nnesha (nnesha):

correct so sinx = 0 and sinx = -1 you can use the unit circle or triangles theorem to find exact solution

OpenStudy (helpjebal):

so since sine = y I'd have to look at the unit circle for coordinates that have a y of -1 or 0?

Nnesha (nnesha):

correct sin theta =y-coordinate

OpenStudy (helpjebal):

thank you!